# Diffusion Generative Recommendations with Continuous Tokens

### v3

@@ Abstract.
Recent advances in generative artificial intelligence, particularly large language models (LLMs), have opened new opportunities for enhancing recommender systems (RecSys). Most existing LLM-based RecSys approaches operate in a discrete space, using vector-quantized tokenizers to align with the inherent discrete nature of language models. However, these quantization methods often result in lossy tokenization and suboptimal learning, primarily due to inaccurate gradient propagation caused by the non-differentiable argmin operation in standard vector quantization. Inspired by the emerging trend of embracing continuous tokens in language models, we propose ContRec, a novel framework that seamlessly integrates continuous tokens into LLM-based RecSys. Specifically, ContRec consists of two key modules: a $\sigma$-VAE Tokenizer, which encodes users/items with continuous tokens; and a Dispersive Diffusion module, which captures implicit user preference. The tokenizer is trained with a continuous Variational Auto-Encoder (VAE) objective, where three effective techniques are adopted to avoid representation collapse. By conditioning on the previously generated tokens of the LLM backbone during user modeling, the Dispersive Diffusion module performs a conditional diffusion process with a novel Dispersive Loss, enabling high-quality user preference generation through next-token diffusion. Finally, ContRec leverages both the textual reasoning output from the LLM and the latent representations produced by the diffusion model for Top-K item retrieval, thereby delivering comprehensive recommendation results. Extensive experiments on four datasets demonstrate that ContRec consistently outperforms both traditional and state-of-the-art LLM-based recommender systems. Our results highlight the potential of continuous tokenization and generative modeling for advancing the next generation of recommender systems. Our code and data are available at https://github.com/Quhaoh233/ContRec.
Web Mining, Recommender Systems, Large Language Models, Diffusion Models.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Web mining is a pivotal research area that focuses on extracting valuable knowledge and patterns from vast and heterogeneous web data. Within this domain, recommender systems (RecSys) serve as crucial tools to tackle information overload and enrich user experiences on various web applications such as e-commerce, job search, and social media platforms (Fan et al., 2019, 2020). The emerging focus on Generative Artificial Intelligence (GenAI), especially Large Language Models (LLMs), has spurred a growing interest in Generative Recommendation, driven by the desire to leverage their exceptional generalization capabilities and adeptness at in-context learning (Zhao et al., 2024; Wang et al., 2025b). These distinct abilities empower LLM-based RecSys to generalize more effectively across new tasks and domains, thereby boosting their overall performance and utility.
@@
Such limitations of discrete learning-based models have garnered attention in language tasks, especially for multimodal modeling (e.g., image, video, and protein) (Wang et al., 2024a, 2025c). Li et al. (2024) suggest that auto-regressive models do not necessarily need to be coupled with discrete and quantized representations for image generation. Fan et al. (2025a) reveal that quantization-based visual models exhibit poor performance improvements when scaling up model size, compared to models operating on continuous tokens. Furthermore, there have been increasing efforts to introduce continuous tokens for learning better representations. For example, Chen et al. (2025) introduce an efficient continuous image tokenizer for language models, which achieves a high compression ratio while enriching the semantic content of the latent space for image reconstruction. These advancements in language models inspire the exploration of continuous tokens as a new tokenization paradigm for LLM-based generative recommendation. To achieve the goal, a few LLM-based recommendation methods have attempted to include continuous tokens. For example, CoLLM (Zhang et al., 2025) and LlaRA (Liao et al., 2024) propose directly projecting collaborative item representations learned from traditional recommendation methods into the token space of language models. However, these methods only continualize the input part of the model by directly employing simple projection layers, e.g., a multi-layer perceptron (MLP), for user/item tokenization, while their output part inherits the LLM backbone’s generator, which is learned for in-vocabulary discrete tokens. Such discrete-continuous gaps in one model might degrade its unification (Sun et al., 2024) and result in suboptimal recommendation performance.
@@
To enable the seamless incorporation of continuous tokens into LLM-based generative recommendation, we propose leveraging the powerful Diffusion Model to continualize user preference modeling and generation, alongside the tokenization of users and items. It is widely recognized that user preference modeling is the core technical part leading to effective recommendation, while this task is highly challenging because user preferences are inherently complex and differences between individual users can be exceptionally subtle. Diffusion Models are particularly promising for addressing this challenge, as they have demonstrated remarkable success in handling continuous tokens across various domains. For instance, DEEM (Luo et al., 2025) demonstrates that diffusion models can serve as the ”eyes” of LLMs for image perception by encoding and decoding images as continuous tokens, while recent advances in protein language models demonstrate diffusion models’ capability in constructing continuous protein structure embeddings (Guo et al., 2024; Wang et al., 2024b). To investigate the potential of continuous tokens and diffusion models in recommendation scenarios, we conduct a preliminary experiment comparing their performance in reconstructing item embeddings on the Amazon-Beauty dataset. As shown in Figure 1, our results are consistent with previous findings in the literature, showing that utilizing diffusion models with continuous representations delivers higher reconstruction accuracy and a smoother loss decline than the discrete counterparts. Detailed experimental configurations and further discussion are provided in Section 4.3.3 and Appendix C.
@@
In this paper, we present a novel LLM-based recommendation framework, ContRec, utilizing Continuous tokens to enhance the capabilities of LLM-based RecSys. With an LLM being the recommender backbone for user modeling with hybrid input, ContRec consists of two core continuous token learning components: a VAE module to learn continuous tokenization for users/items via the reconstruction of their associated features (e.g., IDs, texts, and images) and a novel Disperse Diffusion module to learn continuous-distributed user preference representations based on the LLM-based user modeling. Specifically, our pre-trained VAE model generates continuous tokens for users and items, which are fed into the LLM along with the descriptive text for each recommendation case of user modeling. The Disperse Diffusion module is further applied to take LLM-output user modeling results as conditions to generate continuous user preference representations. Finally, we design a hybrid retrieval mechanism that integrates the explicit LLM-output prediction text and the implicit Diffusion-enhanced user preference representation to output a ranking list of items for precise recommendations. Overall, our model enables the modeling of continuous-distributed users and items as well as enhanced user preferences, thereby greatly boosting the recommendation performance.
@@ 2. Preliminary
Diffusion Models. We provide a brief overview of diffusion models, focusing on the representative formulation known as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020). DDPM leverages the chain rule of probability and the Markov property to model the underlying data-generating distribution. In the forward process, DDPM gradually adds Gaussian noise to the original input $\mathbf{x}^{0}$ according to a variance schedule $\beta_{t}$:
@@
where the initial state $\mathbf{x}^{T}$ is sampled from standard Gaussian noise $\mathbf{x}^{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Here, $\eta(\mathbf{x}^{t},t)$ denotes a time-dependent mean of the Gaussian distribution used to sample the previous timestep’s data (often initialized as untrained constants), and $\xi(\mathbf{x}^{t},t)$ determines the variance (or covariance) of the Gaussian distribution used in the reverse denoising process. Given the remarkable success of diffusion models in handling continuous data such as images and protein structures, an increasing number of studies have explored integrating diffusion models with large language models (Sun et al., 2024; Wang et al., 2024b, a). This integration enables these models to achieve enhanced expressive capabilities in multimodal modeling and high-quality generation based on implicit representations.
@@ 3.1. Overview of ContRec
In order to leverage continuous tokens into an LLM-based recommender system to enhance the model’s capability for user, item, and implicit user preference representation, we propose a novel framework: ContRec. As illustrated in Figure 2, our ContRec consists of two core innovative modules, the $\sigma$-VAE Tokenizer and Disperse Diffusion. The $\sigma$-VAE module performs robust and non-quantized tokenization on users and items based on VAE (Kingma and Welling, 2014; Rezende et al., 2014) to obtain continuous tokens. We further incorporate several simple yet effective techniques to prevent the collapse of latent representations during the training process. With continuous tokens obtained, ContRec processes interleaved text and recommendation tokens as input, leveraging the powerful reasoning capabilities of LLMs to model user preferences. On one hand, the hidden states produced by the LLM backbone are decoded into textual descriptions of user preferences via auto-regressive modeling. On the other hand, these hidden states are also provided as conditions to a specially designed diffusion model, which generates continuous representations of user preferences. Within this process, we introduce a novel Disperse Loss (Wang and He, 2025), which simulates contrastive learning for recommendation without the need for negative sampling. Finally, we employ a hybrid retrieval process that integrates the outputs from both the LLM backbone and the diffusion model, enabling the effective generation of top-K recommendations. Due to space constraints, we present only the core formulas of the key components of the proposed method in this section; supplementary information and detailed derivations are provided in Appendix B.
@@ 3.2. $\sigma$-VAE Tokenizer
The proposed tokenizer indexes users and items using latent vectors. It is built upon the VAE framework (Kingma and Welling, 2014), which encodes input data into a latent space and subsequently decodes it back to the original space. Maintaining a larger variance in the latent representations improves the robustness of the VAE tokenizer to exposure bias between tokens during inference, while standard VAEs often suffer from channel variance collapse, adversely affecting autoregressive modeling (Razavi et al., 2019). To address this issue, we introduce three simple yet effective techniques to enhance our Tokenizer: (i) a masking operation based on the Bernoulli distribution to enhance robustness, (ii) a K-way encoder with parallel encoding channels to allow implicit encoding (Qu et al., 2025), and (iii) a Gaussian kernel applied to each channel to prevent variance collapse (Sun et al., 2024). The reconstruction pass is computed as:
@@ 3.4. Dispersive Diffusion
Denoising diffusion models have recently demonstrated robust capabilities in modeling complex data distributions and learning continuous representation spaces for various modalities, including images (Li et al., 2024), protein structures (Wang et al., 2024b), and videos (Sun et al., 2024). Transferring this success to LLM-based recommender systems is a promising direction to help with the delicate user preference modeling (Ho et al., 2020; Nichol and Dhariwal, 2021). To investigate the feasibility of applying diffusion models to LLM-based recommender systems, we conduct a mathematical analysis that proves the integration of continuous representations in the LLM backbone (Li et al., 2024). Specifically, language models formulate the generation recommendation problem as “next token prediction” in an autoregressive manner as follows:
@@
(18) $\mathcal{L}_{\text{contrast}}=-\log\frac{\exp(-\mathcal{D}(\hat{\mathbf{y}}_{i},{\mathbf{y}}_{i})/\tau)}{\sum_{m}\exp(-\mathcal{D}(\hat{\mathbf{y}}_{i},\mathbf{y}_{m})/\tau)},$
where $(\hat{\mathbf{y}_{i}},\mathbf{y}_{m})$ represents sample pairs, including both the positive pair and all negative pairs. $\mathcal{D}(\cdot)$ is the squared $\ell_{2}$ distance, and $\tau$ is a hyper-parameter known as the diffusion temperature that controls the learning speed of the diffusion process.
@@
(19) $\mathcal{L}_{\text{disp}}=\log\mathbb{E}_{i,b}[\exp(-\mathcal{D}(\mathbf{h}_{i},\mathbf{h}_{b})/\tau)],$
@@ 3.5.2. Inference
For the Diffusion process (Ho and Salimans, [n. d.]) during inference, we introduce a hyperparameter $\omega$ that controls the strength of conditioning to modulate the effect of the guidance signal $\mathbf{c}_{i}$. We use $\omega$ to modify the noise prediction in Eq. (17). While a higher value of $\omega$ can enhance personalized guidance, it may also reduce the generalization ability of the diffusion process, potentially deteriorating the quality of the generated oracle item. Mathematically, the modified noise prediction process can be formulated as follows:
@@ 4.1.1. Datasets
We conducted a series of comprehensive experiments across four benchmark datasets to evaluate the effectiveness of the proposed ContRec and core modules in it. First, LastFM, sourced from the HetRec 2011111https://grouplens.org/datasets/hetrec-2011/ repository, encompasses music artist listening data from users within the Last.fm online music platform. Moreover, the ML1M dataset offers a collection of movie ratings made by MovieLens [^2] https://grouplens.org/datasets/movielens/1m/ users. The last two datasets—Beauty and Video Games—are sourced from the Amazon Review Data [^3] https://nijianmo.github.io/amazon/ repository. These datasets have undergone a reduction process to derive their 5-core versions, ensuring that each user and item retains a minimum of five reviews. Table 1 details the characteristics of these datasets, with the maximum item sequence length capped at 20 to align with a 1024-token input constraint. In our cases, each item’s title, ID, brand, and category are used, while only the ID is considered on the user side. Moreover, our dataset division follows the Split-by-Timepoint policy (Ji et al., 2023), in which all interactions occurring after a predefined time point are designated as test instances, while those before the time point are used for training. To ensure that all items appear at least once in the training corpus, we split the sequential recommendation data in chronological order at the 90th and 95th percentiles to create the training, validation, and testing sets.
@@ 4.1.2. Compared Models
Here, we compare our approach with four representative sequential recommendation methods and five recently developed LLM-based RecSys. The four sequential recommendation methods are GRU4Rec (Hidasi and Karatzoglou, 2018), SASRec (Kang and McAuley, 2018), SSD4Rec (Qu et al., 2024b), and DreamRec (Yang et al., 2023), with DreamRec being a diffusion-based method. The five LLM-based models are: (i) P5 (Geng et al., 2022) reformats recommendation tasks into a text-to-text format and finetunes a small pre-trained language model (i.e., T5) to capture more profound semantics for recommendation. (ii) CoLLM (Zhang et al., 2025) incorporates collaborative representations of items into LLMs using a projection layer. (iii) TIGER (Rajput et al., 2023) employs residual vector quantization to distill large volumes of text into a compact set of semantic identifiers, and then leverages a Transformer-based architecture trained on these semantic ID sequences to perform sequential recommendation. (iv) LLaRA (Liao et al., 2024) uses a projector to align item embeddings with textual item information and implements a curriculum learning approach for finetuning. (v) TokenRec (Qu et al., 2025) proposes a novel quantized tokenizer to encode users and items along with their collaborative knowledge into discrete tokens that are compatible with LLMs.

### v4

@@ Abstract.
Recent advances in generative artificial intelligence, particularly large language models (LLMs), have opened new opportunities for enhancing recommender systems (RecSys). Most existing LLM-based RecSys approaches operate in a discrete space, using vector-quantized tokenizers to align with the inherent discrete nature of language models. However, these quantization methods often result in lossy tokenization and suboptimal learning, primarily due to inaccurate gradient propagation caused by the non-differentiable argmin operation in standard vector quantization. Inspired by the emerging trend of embracing continuous tokens in language models, we propose ContRec, a novel framework that seamlessly integrates continuous tokens into LLM-based RecSys. Specifically, ContRec consists of two key modules: a $\sigma$-VAE Tokenizer, which encodes users/items with continuous tokens; and a Dispersive Diffusion module, which captures implicit user preference. The tokenizer is trained with a continuous Variational Auto-Encoder (VAE) objective, where three effective techniques are adopted to avoid representation collapse. By conditioning on the previously generated tokens of the LLM backbone during user modeling, the Dispersive Diffusion module performs a conditional diffusion process with a novel Dispersive Loss, enabling high-quality user preference generation through next-token diffusion. Finally, ContRec leverages both the textual reasoning output from the LLM and the latent representations produced by the diffusion model for Top-K item retrieval, thereby delivering comprehensive recommendation results. Extensive experiments on four datasets demonstrate that ContRec consistently outperforms both traditional and state-of-the-art LLM-based recommender systems. Our results highlight the potential of continuous tokenization and generative modeling for advancing the next generation of recommender systems.
Recommender Systems, Large Language Models, Diffusion Models.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792428isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Web mining
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.17285302.
@@ 1. Introduction
Web mining is a pivotal research area that focuses on extracting valuable knowledge and patterns from vast and heterogeneous web data. Within this domain, recommender systems (RecSys) serve as crucial tools to tackle information overload and enrich user experiences on various web applications such as e-commerce, job search, and social media platforms (Wang et al., 2020; Fan et al., 2019, 2020). The emerging focus on Generative Artificial Intelligence (GenAI), especially Large Language Models (LLMs), has spurred a growing interest in Generative Recommendation, driven by the desire to leverage their exceptional generalization capabilities and adeptness at in-context learning (Zhao et al., 2024; Wang et al., 2025b). These distinct abilities empower LLM-based RecSys to generalize more effectively across new tasks and domains.
@@
Such limitations of discrete learning-based models have garnered attention in language tasks, especially for multimodal modeling (e.g., image, video, and protein) (Wang et al., 2024, 2025c). For example, Fan et al. (2025a) reveal that quantization-based visual models exhibit poor performance improvements when scaling up model size, compared to models operating on continuous tokens. Furthermore, there have been increasing efforts to introduce continuous tokens for learning better representations (Li et al., 2024). For example, Chen et al. (2025) introduce an efficient continuous image tokenizer for language models, which achieves a high compression ratio while enriching the semantic content of the latent space for image reconstruction. These advancements in language models inspire the exploration of continuous tokens as a new tokenization paradigm for LLM-based generative recommendation. To achieve the goal, a few LLM-based recommendation methods have attempted to include continuous tokens. For example, CoLLM (Zhang et al., 2025a) and LlaRA (Liao et al., 2024) propose directly projecting collaborative item representations learned from traditional recommendation methods into the token space of language models. However, these methods only continualize the input part of the model by directly employing simple projection layers, e.g., a multi-layer perceptron (MLP), for user/item tokenization, while their output part inherits the LLM backbone’s generator, which is learned for in-vocabulary discrete tokens. Such discrete-continuous gaps in one model might degrade its unification (Sun et al., 2024) and result in suboptimal recommendation performance.
@@
To enable the seamless incorporation of continuous tokens into LLM-based generative recommendation, we propose leveraging the powerful Diffusion Model to continualize user preference modeling and generation, alongside the tokenization of users and items. It is widely recognized that user preference modeling is the core technical part leading to effective recommendation, while this task is highly challenging because user preferences are inherently complex and differences between individual users can be exceptionally subtle. Diffusion Models are particularly promising for addressing this challenge, as they have demonstrated remarkable success in handling continuous tokens across various domains. For instance, DEEM (Luo et al., 2025) demonstrates that diffusion models can serve as the ”eyes” of LLMs for image perception by encoding and decoding images as continuous tokens, while recent advances in protein language models demonstrate diffusion models’ capability in constructing continuous protein structure embeddings (Guo et al., 2024; Wang et al., 2025d). To validate this insight in recommendation scenarios, we conduct a preliminary item embedding reconstruction experiment on the Amazon-Beauty dataset. As shown in Figure 1, our results are consistent with previous findings in the literature, showing that utilizing diffusion models with continuous representations delivers higher reconstruction accuracy and a smoother loss decline than the discrete counterparts. Detailed information is provided in Section 4.3.3 and Appendix C.
@@
In this paper, we present a novel LLM-based recommendation framework, ContRec, utilizing Continuous tokens to enhance the capabilities of LLM-based RecSys. With an LLM being the recommender backbone for user modeling with hybrid input, ContRec consists of two core continuous token learning components: a VAE module to learn continuous tokenization for users/items via the reconstruction of their associated features (e.g., IDs, texts, and images) and a novel Disperse Diffusion module to learn continuous-distributed user preference representations based on the LLM-based user modeling. Additionally, we design a hybrid retrieval mechanism that integrates the explicit LLM-output prediction text and the implicit diffusion-enhanced user preference representation to output a ranking list of items for precise recommendations.
@@ 2. Preliminary
Diffusion Models. We provide a brief overview of diffusion models, focusing on the representative formulation known as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020). In the forward process, DDPM gradually adds Gaussian noise to the original input $\mathbf{x}^{0}$ according to a variance schedule $\beta_{t}$:
@@
where the initial state $\mathbf{x}^{T}$ is sampled from standard Gaussian noise $\mathbf{x}^{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Here, $\eta(\mathbf{x}^{t},t)$ denotes a time-dependent mean of the Gaussian distribution used to sample the previous timestep’s data (often initialized as untrained constants), and $\xi(\mathbf{x}^{t},t)$ determines the variance (or covariance) of the Gaussian distribution used in the reverse denoising process. Given the remarkable success of diffusion models in handling continuous data such as images and protein structures, an increasing number of studies have explored integrating diffusion models with LLMs (Sun et al., 2024; Wang et al., 2025d, 2024).
@@ 3.1. Overview of ContRec
In order to leverage continuous tokens into an LLM-based recommender system to enhance the model’s capability for user, item, and implicit user preference representation, we propose a novel framework: ContRec. As illustrated in Figure 2, our ContRec consists of two core innovative modules, the $\sigma$-VAE Tokenizer and Disperse Diffusion. Due to space constraints, we present only the core formulas of the key components of the proposed method in this section; supplementary information and detailed derivations are provided in Appendix B.
@@ 3.2. $\sigma$-VAE Tokenizer
The proposed tokenizer is built upon the VAE framework (Kingma and Welling, 2014; Rezende et al., 2014), which encodes input data into a latent space and subsequently decodes it back to the original space. Maintaining a larger variance in the latent representations improves the robustness of the VAE tokenizer to exposure bias between tokens during inference, while standard VAEs often suffer from channel variance collapse, adversely affecting autoregressive modeling (Razavi et al., 2019). To address this issue, we introduce three simple yet effective techniques to enhance our Tokenizer: (i) a masking operation based on the Bernoulli distribution to enhance robustness, (ii) a K-way encoder with parallel encoding channels to allow implicit encoding (Qu et al., 2025), and (iii) a Gaussian kernel applied to each channel to prevent variance collapse (Sun et al., 2024). The reconstruction pass is computed as:
@@ 3.4. Dispersive Diffusion
Denoising diffusion models have recently demonstrated robust capabilities in modeling complex data distributions and learning continuous representation spaces for various modalities, including images (Li et al., 2024), protein structures (Wang et al., 2025d), and videos (Sun et al., 2024). Transferring this success to LLM-based recommender systems is a promising direction to help with the delicate user preference modeling (Ho et al., 2020; Nichol and Dhariwal, 2021). To investigate the feasibility of applying diffusion models to LLM-based recommender systems, we conduct a mathematical analysis that proves the integration of continuous representations in the LLM backbone (Li et al., 2024). Specifically, language models formulate the generation recommendation problem as “next token prediction” in an autoregressive manner as follows:
@@
(18) $\mathcal{L}_{\text{contrast}}=-\log\frac{\exp(-\mathcal{D}(\hat{\mathbf{y}}_{i},{\mathbf{y}}_{i})/\iota)}{\sum_{m}\exp(-\mathcal{D}(\hat{\mathbf{y}}_{i},\mathbf{y}_{m})/\iota)},$
where $(\hat{\mathbf{y}_{i}},\mathbf{y}_{m})$ represents sample pairs, including both the positive pair and all negative pairs. $\mathcal{D}(\cdot)$ is the squared $\ell_{2}$ distance, and $\iota$ is a temperature parameter for contrastive learning, which also controls the learning speed of the diffusion process in our cases.
@@
(19) $\mathcal{L}_{\text{disp}}=\log\mathbb{E}_{i,b}[\exp(-\mathcal{D}(\mathbf{h}_{i},\mathbf{h}_{b})/\iota)],$
@@ 3.5.2. Inference
For the Diffusion process (Ho and Salimans, [n. d.]) during inference, we introduce a hyperparameter $\omega$ that controls the strength of conditioning to modulate the effect of the guidance signal $\mathbf{c}_{i}$. We use $\omega$ to modify the noise prediction in Eq. (17). While a higher value of $\omega$ can enhance personalized guidance, it may also reduce the generalization ability of the diffusion process, potentially deteriorating the quality of the generated oracle item. Mahematically, the modified noise prediction process can be formulated as follows:
@@ 4.1.1. Datasets
We conducted a series of comprehensive experiments across four benchmark datasets to evaluate the effectiveness of the proposed ContRec and core modules in it. First, LastFM, sourced from the HetRec 2011111https://grouplens.org/datasets/hetrec-2011/ repository, encompasses music artist listening data from users within the Last.fm online music platform. Moreover, the ML1M dataset offers a collection of movie ratings made by MovieLens [^2] https://grouplens.org/datasets/movielens/1m/ users. The last two datasets—Beauty and Video Games—are sourced from the Amazon Review Data [^3] https://nijianmo.github.io/amazon/ repository. These datasets have undergone a reduction process to derive their 5-core versions, ensuring that each user and item retains a minimum of five reviews. Table 1 details the characteristics of these datasets, with the maximum item sequence length capped at 20 to align with a 1024-token input constraint. In our cases, each item’s title, ID, brand, and category are used, while only the ID is considered on the user side. Moreover, our dataset division follows the Split-by-Timepoint policy (Ji et al., 2023), in which all interactions occurring after a predefined time point are designated as test instances, while those before the time point are used for training. To ensure that all items appear at least once in the training corpus, we split the recommendation data in chronological order at the 90th and 95th percentiles to create the training, validation, and testing sets.
@@ 4.1.2. Compared Models
Here, we compare our approach with four representative sequential recommendation methods and five recently developed LLM-based RecSys. The four sequential recommendation methods are GRU4Rec (Hidasi and Karatzoglou, 2018), SASRec (Kang and McAuley, 2018), SSD4Rec (Zhang et al., 2025b), and DreamRec (Yang et al., 2023), with DreamRec being a diffusion-based method. The five LLM-based models are: (i) P5 (Geng et al., 2022) reformats recommendation tasks into a text-to-text format and finetunes a small pre-trained language model (i.e., T5) to capture more profound semantics for recommendation. (ii) CoLLM (Zhang et al., 2025a) incorporates collaborative representations of items into LLMs using a projection layer. (iii) TIGER (Rajput et al., 2023) employs residual vector quantization to distill large volumes of text into a compact set of semantic identifiers, and then leverages a Transformer-based architecture trained on these semantic ID sequences to perform sequential recommendation. (iv) LLaRA (Liao et al., 2024) uses a projector to align item embeddings with textual item information and implements a curriculum learning approach for finetuning. (v) TokenRec (Qu et al., 2025) proposes a novel quantized tokenizer to encode users and items along with their collaborative knowledge into discrete tokens that are compatible with LLMs.